OpenStudy (anonymous):

hELP :) ? -10 = -2 + 1/2 d

OpenStudy (anonymous):

d=-16

OpenStudy (anonymous):

-10 = -2 + 1/2 d -10+2=-2 +(1/2)(d)+2 <--- to put x on one side and numbers at the other side -8=(1/2)(d) (-8)(2)=(1/2)(d)(2) <-- to remove 1/2 -16=d therefore, d=-16